OpenStudy (anonymous):

(a) Ten inches of snow is equivalent to about one inch of rain.1 Write an equation for the amount of precipitation, measured in inches of

OpenStudy (anonymous):

So you're looking for the amount of rain given the amount of snow. First let's label y as the amount of rain and x as the amount of snow. To get the equation use a ratio. (1 in rain)/(10 in snow) = y/x Cross multiply to get the final equation. \[y = (1/10) x\] x= amount of snow in inches y= amount of rain in inches